eC https:/session.masteringphysics.com/my KHomework Assignment 1 (Chapter 2) Problem C2-06 Constants A car travels north at 29.700 m/s. The driver applies the brakes and slows the car to 14.000 m/s for 140.40 m Assume the motion is at uniform acceleration.
1)What is the magnitude of the acceleration in m/s^2?
2)What is the direction of the acceleration? (N,S,E,W).
3)If the brakes are applied until the care comes tia stop, what is the total distance covered?
4) how much time elapsed between the instant when the brakes were applied and when the car had traveled a distance of d/2?
Answer #1

Steps

1) use equation of motion

v^2 = u^2 + 2*a*s , to get accelaration

2 ) direction of retardation accelaration is always opposite to the direction of motion

3) Again use equation of step 1 to get total distance covered

4) Again use ist step equation to get velocity at a distance d/2

Now use equation v= u + a* t

to get time elapsed .velocity , after 6、ake O m PART -1 2. 1 ナ2a-S a c 2.9933 ms magnitule a 2.4933 m/sART -2 As ite accelaratnhetandenm ei nature an ae osite Che direetion of motcon A -- 2.443 m/s 2 2 vel 1g6.53 m atance havelleo res :1803 m

PART-4 80.53 m 2 90. 24Sm 2 at the enstanl when 7u car 2. 2. 3 5 see elabstof befveen he uifanl